Need help finding the median for the given sample data.
{ -8, -5, 5, 11}
Please!

Found 2 solutions by solver91311, palanisamy:
Answer by solver91311(24713)



When there are an odd number of data points, the median is the middle point when the data points are put into numerical order. When there are an even number of data points, again put into numerical order, the median is the numerical average of the two in the middle. The numerical average is the same as the mean: add the two center data points and divide the result by 2.

John


Answer by palanisamy(496)

The given data is -8,-5,5,11
Median = (-5+5)/2 = 0/2 = 0